# Comparing strange and non-strange quark stars within resummed QCD at NLO

### v1

@@ Abstract
We employ the renormalization group optimized perturbation theory (RGOPT) resummation method to evaluate the equation of state (EoS) for strange ($N_{f}=2+1$) and non-strange ($N_{f}=2$) cold quark matter at NLO. This allows us to obtain the mass-radius relation for pure quark stars and compare the results with the predictions from perturbative QCD (pQCD) at NNLO. Choosing the renormalization scale to generate maximum star masses of order $M=2-2.6M_{\odot}$, we show that the RGOPT can produce mass-radius curves compatible with the masses and radii of some recently observed pulsars, regardless of their strangeness content. The scale values required to produce the desired maximum masses are higher in the strange scenario since the EoS is softer in this case. The possible reasons for such behavior are discussed. Our results also show that, as expected, the RGOPT predictions for the relevant observables are less sensitive to scale variations than those furnished by pQCD.
@@ I Introduction
Nevertheless, it would be legitimate to argue that such a result could be an artifact of the model approximation adopted in Ref. [20]. After all, when compared to QCD, the (effective) NJL model has some limitations which include the absence of asymptotic freedom and non-renormalizability. On the other hand, the determination of the QCD EoS for cold and compressed matter, from ab initio evaluations, presents further issues which have not yet been circumvented. For instance, within this regime lattice QCD (LQCD) simulations are still plagued by the infamous sign problem [33, 34] while perturbative QCD (pQCD) applications are only reliable at extremely high baryon densities, of order $n_{B}\sim 40n_{0}$ [35, 36, 37, 38] ($n_{0}=0.16\,{\rm fm}^{-3}$), where asymptotic freedom allows for weak coupling expansions. Regarding pQCD it is important to mention that when the modified minimal subtraction ($\overline{\rm MS}$) renormalization scale ($\Lambda$) is taken at the conventional “central” value, $32(\mu_{u}+\mu_{d}+\mu_{s})/$ 3, with $\mu_{f}$ the quark flavor chemical potentials, the NLO pQCD predicts QS masses above $2M_{\odot}$ [39] and below $2M_{\odot}$ at NNLO [40, 41, 35, 42, 43].
@@
Although the cold dense pQCD NNLO pressure is formally perturbatively renormalization group (RG) invariant, the higher order residual scale dependence of pQCD results [35, 36, 37] is still rather sizeable, conventionally taken to lie between half and twice the central scale choice value. Quite recently, pQCD calculations reached NNNLO [44, 45, 46, 47] within the massless quark approximation, and all orders in the leading soft logarithms approximation [48], producing in such cases significant residual scale dependence reductions. However, accounting for the strange quark mass effects, sizeable at moderate and low $\mu$ values, is very involved and presently limited to NNLO, thus impiging on scale uncertainties reductions. In order to tackle the scale dependence problem an alternative method, which better incorporates RG properties within (variationally) optimized perturbative evaluations has been proposed [49]. This renormalization group optimized perturbation theory (RGOPT) technique has been successfully used, at vanishing temperatures and densities, to calculate quantities such as the QCD scale $\Lambda_{\overline{\rm MS}}$ and strong coupling $\alpha_{s}$ [49], and the quark condensate up to four and five loop orders [50, 51]. Later, the RGOPT has been applied to scalar field theories at finite temperatures [52, 53, 54, 55], indicating that the method greatly improves the scale dependence in thermal theories, when compared to screened perturbation theory (SPT) [56, 57] and hard thermal loop perturbation theory (HTLpt) [58, 57, 59] based on the HTL effective theory [60]. The technique was also employed to describe cold and dense QCD [61], as well as the quark sector of QCD at finite temperatures and baryonic densities [62, 63]. Once again, the RGOPT predictions display a residual scale dependence which is smaller than those furnished by pQCD and HTLpt. Recently, some of the present authors have used the RGOPT EoS for QCD in the chiral limit ($m_{u}=m_{d}\equiv 0$) to describe non-strange quark stars (NSQS) obtaining encouraging results at NLO [32]. In particular, when the renormalization scale is appropriately chosen, the RGOPT resummation of the perturbative series is able to describe several observational data, specially from pulsars PSR J0030+0451 [64, 65], PSR J0740+6620 [66, 67, 68] and the light compact object identified in the middle of a supernova remnant, HESS J1731-347 [69].
@@
Here, our major goal is to investigate how the inclusion of strangeness affects the $N_{f}=2$ results obtained in Ref. [32]. In the present $N_{f}=2+1$ case, the introduction of a finite (strange) quark mass ($m_{s}$) modifies the RG operator introducing an extra scale dependence, through the running of $m_{s}(\Lambda)$, which needs to be taken into account by the resummation procedure. This is accordingly a nontrivial extension with respect to the determination of the RGOPT EoS for massless quarks at $N_{f}=3$ [61, 62, 63] and $N_{f}=2$ [32]. Following the usual procedure, here we will also choose the scale to be density dependent so that it can be written in terms of the baryonic chemical potential [^1] In this situation the pressure needs to be modified to ensure thermodynamic consistency. ($\mu_{B}$) as $\Lambda=X\mu_{B}/3$, where $X$ represents a free parameter that can be adjusted in order to reproduce the desired maximum mass value for a given family of stars. Then, the QS properties determined from the novel $N_{f}=2+1$ RGOPT EoS will be compared with those previously obtained with the same method when $N_{f}=2$ [32] as well as with those furnished by $N_{f}=2+1$ pQCD at NNLO [35, 36], for the corresponding flavor species. Since $\alpha_{s}$ runs with $\Lambda=X\mu_{B}/3$ the value of $X$ needed to produce a massive quark star will enable us to compare the coupling strength associated with each different number of flavors. This comparison will then allow us to draw further conclusions about the reliability of the distinct EoS provided by the RGOPT at NLO.
@@ II Perturbative QCD pressure for $N_{f}=2+1$ at NLO
$\begin{split}P^{\rm{PT}}_{f}(m_{f},\mu)=&-N_{c}\frac{m_{f}^{4}}{8\pi^{2}}\left(\frac{3}{4}-L_{m}\right)+\Theta(\mu^{2}-m_{f}^{2})\,\frac{N_{c}}{12\pi^{2}}\left[\mu p_{F}\left(\mu^{2}-\frac{5}{2}m_{f}^{2}\right)+\frac{3}{2}m_{f}^{4}\ln(\frac{\mu+p_{F}}{m_{f}})\right]\\ &-\frac{g^{2}d_{A}}{4\left(2\pi\right)^{4}}m_{f}^{4}\left(3L_{m}^{2}-4L_{m}+\frac{9}{4}\right)-\Theta(\mu^{2}-m_{f}^{2})\,\frac{g^{2}d_{A}}{4\left(2\pi\right)^{4}}\left\{3\left[m_{f}^{2}\ln(\frac{\mu+p_{F}}{m_{f}})-\mu p_{F}\right]^{2}-2p_{F}^{4}\right\}\\ &-\Theta(\mu^{2}-m_{f}^{2})\,\frac{g^{2}d_{A}}{4\left(2\pi\right)^{4}}m_{f}^{2}\left(4-6L_{m}\right)\left[\mu p_{F}-m_{f}^{2}\ln(\frac{\mu+p_{F}}{m_{f}})\right]\,,\end{split}$ (1)
@@ III RGOPT resummation setup
At this point, some important additional remarks are in order: First, note that although these RG properties were originally obtained from vacuum ($\mu=T=0$) contributions, they still hold when thermal and/or in-medium effects are considered [53, 52, 54, 62, 63]). Also, according to the original prescription, $a$ is fixed once for all at LO, and the same value in Eq. (15) is used at successive perturbative orders of the (modified) perturbative $\delta$-expansion. Importantly, this guarantees [49] in particular that at higher orders there is always one solution $\overline{\eta}(g)$ compatible with asymptotic freedom behavior for $g\to 0$. Finally, note that even when treating massive theories as we do here, it is compelling to adopt this universal constant value for $a$ in Eq. (15), in order to avoid otherwise a clumsy $a(g,m,...)$ dependent exponent, furthermore incompatible with standard renormalization. We emphasize that although other slightly different variational prescriptions could be thought of, the latter one provides a sensible way of comparing successive perturbative orders with the same prescription. This is further justified a posteriori since the strange quark mass ($m_{s}$) which is relevant here remains a rather moderate perturbation [^4] As we will see in the sequel, $m_{s}\ll\overline{\eta}\simeq{\cal O}(g\,\mu)\ll\mu$ roughly holds, at least as long as the QCD coupling, $g^{2}(\Lambda)$, remains moderate (i.e. for renormalization scales $\Lambda\sim{\cal O}(\mu)$ with $\mu$ not too small). with respect to the dressed mass $\eta(g^{2},\mu)$ determined from Eq. (17).
@@ III.1 Optimization
After applying this RSC to the original NLO pressure, one must re-expand perturbatively the latter to NLO, producing an extra term $-4g^{2}\,m_{f}^{4}s_{0}B_{2}$ in the resulting pressure. As a by-product, after the RGOPT modifications, similar contributions with $m_{f}\to\eta+m_{s}$ and $m_{f}\to\eta$ are induced respectively in the first and secont term on the right hand side of Eq. (21). Although the RSC parameters are arbitrary to start with, one seeks a sensible prescription to recover real solutions while maintaining $B_{2}$ reasonably close to $\overline{\rm MS}$ results to justify a posteriori such a perturbative RSC. As Ref. [49] prescribes, $B_{2}$ is accordingly fixed (uniquely) upon requiring the nearest “contact” of the two curves parametrizing the MOP Eq. (17) and RG Eq. (16), considered as functions $f_{\rm{MOP}}(\eta,g^{2})$, $f_{\rm{RG}}(\eta,g^{2})$ respectively. Algebraically this corresponds to solving
@@
Fig. 1 compares the optimized mass parameter, $\overline{\eta}$, as well as the normalized pressure as functions of the baryon chemical potential representing symmetric matter, $\mu_{B}=3\mu$, for $N_{f}=3$ and $N_{f}=2+1$. Following the standard prescription, we have considered three different values of the renormalization scale, $\Lambda=X\mu_{B}/3$, with $X=1,2,4$, to quantify the residual scale dependence. Regarding the variational masses, the figure suggests that they share a similar behavior, increasing linearly with $\mu_{B}$ roughly as $\overline{\eta}\sim{\cal O}(g\,\mu)$, although $\overline{\eta}$ actually entails an all-order $g^{2}$-dependence from RG resummation. Importantly, the fact that $\overline{\eta}$ accordingly behaves as a screening mass for moderate coupling $g$, also guarantees that the RSC Eq. (23) modifications induced to the pressure, $\propto-B_{2}s_{0}g^{2}\overline{\eta}^{4}$, are of formally higher orders ${\cal O}(g^{6})$. For the same $\mu_{B}$ value, and comparing with both $N_{f}=2$ and $N_{f}=3$ cases, the presence of a massive quark ($N_{f}=2+1$) generates a lower pressure. This can be explained by first noting that in the basic NLO pressure Eq. (1) or (2), the $m_{f}\neq 0$ contributions are always negative as compared to the equivalent massless contributions (at least as long as $m_{f}\ll\mu$), and this feature remains true for the modified NLO expression Eq. (19). Then since $\overline{\eta}+m_{s}>\overline{\eta}$, the lowering of the pressure is obviously more pronounced for $m_{s}\neq 0$ in the RGOPT case [^7] In contrast, the RSC modification $\sim-B_{2}s_{0}g^{2}\overline{\eta}^{4}$ gives a positive extra contribution to the pressure, as $s_{0}>0$ in Eq. (6) and we always obtain $B_{2}<0$. But this contribution is subdominant compared to the ”direct” $\overline{\eta}$ dependence entering other contributions.. A lower pressure tends to enhance the remnant scale dependence, therefore slightly counteracting the scale dependence improvement obtained from RGOPT, although the figure indicates that the observed scale dependence is not so much sensitive to the strange quark mass value.
@@ IV Chemical equilibrium and thermodynamic consistency
must be observed. The relevant baryon chemical potential describing each different number of flavors is defined as $\mu_{B}=\mu_{u}+2\mu_{d}$ (recall that $\mu_{d}=\mu_{s}$). Seeking a fair comparison, the renormalization scale given by $\Lambda=X\,\mu_{B}/3$ is chosen in all cases. To obtain the pQCD results at $N_{f}=2$, we shall use the NNLO result for massless quarks given by [40, 41, 77],
@@ V Numerical Results
In this section, we present the numerical results furnished by the RGOPT and pQCD for NSQM ($N_{f}=2$) as well as SQM ($N_{f}=2+1$) considering $\beta$-equilibrium, Eqs. (27) and (29), and charge neutrality, Eqs. (30) and (28). The EoS corresponding to each case is used to obtain the mass-radii relation describing non-strange and strange quark stars.
@@
In Fig. 2, the electron chemical potential predicted by the NLO RGOPT is plotted (on the left panel) as a function of $\mu_{B}$ for $N_{f}=2+1$ (solid lines) and $N_{f}=2$ (dashed lines). In the case of SQM, the appearance of kinks in the curves signals the onset of strange quarks which induces a strong reduction of the number of electrons decreasing the electron chemical potential ($\mu_{e}\to 0$, at high-$\mu_{B}$). Therefore, in this limit, the densities of the three quark flavors become similar. In other words, the system favors the replacement of the negative electric charge of electrons by the one of $s$-quarks. On the other hand, in the case of NSQM, as $\mu_{B}$ takes on higher values, the electron production increases to compensate for the increase in the $u$ quark number density, while maintaining charge neutrality. Note that the $u$ quark has a smaller Fermi momentum than the $d$ quark and, therefore, an increase of the number density is favored.
@@
Figure 3: Thermodynamic consistent NLO RGOPT pressure as function of $\mu_{B}$ for $N_{f}=2+1$ (continuous lines) and $N_{f}=2$ (dashed lines). For comparison the dot-dashed and dotted lines respectively represent the $N_{f}=2+1$ and $N_{f}=2$ NNLO pQCD pressure. For all curves the upper ones correspond to $X=4$, the lower ones to $X=1$ while the central scale, $X=2$, is represented by the inner curves. All pressures have been normalized by the Fermi-Dirac value $P_{\rm FD}$ for free massless quarks.
@@
In Fig. 3, we show the normalized pressures furnished by the NLO RGOPT and NNLO pQCD for both SQM and NSQM, thermodynamically consistent after applying Eq. (34). The first thing to note is that, within both approximations, the introduction of a finite strange quark mass decreases the scale dependence (this effect is much more noticeable in the case of pQCD). While those features result from involved combined effects from the running coupling, running mass, and extra thermodynamic consistency contributions, this behavior is explained partly from the fact that since $b_{0}(N_{f}=3)<b_{0}(N_{f}=2)$, the running coupling $g^{2}(\Lambda)\sim[2b_{0}\ln(\Lambda/\Lambda_{QCD})\,]^{-1}$ has somewhat smaller scale variation for $N_{f}=3$ than for $N_{f}=2$. The latter effect is however partly compensated by the sizable massive contributions $\overline{\eta}\neq 0$, lowering the pressure (as explained in the last paragraph before Sec. III), which rather tends to increase the remnant scale dependence. In addition, the figure suggests that the NLO RGOPT produces EoSs that converge slower to the Fermi-Dirac limit (free massless quarks) than its NNLO pQCD counterpart. This happens because the former procedure produces a non-trivial medium-dressed mass $\overline{\eta}(g,\mu)$, so that the effective masses entering the quark propagators are $m_{s}+\overline{\eta}(g,\mu)$ (strange sector) and $\overline{\eta}(g,\mu)$ (non-strange sector). Accordingly, at $N_{f}=2$ the RGOPT quark propagator already involves a non-trivial effective quark mass while, at $N_{f}=2+1$, the effect of including the $s$-quark physical (current) mass is more diluted. A similar behavior was initially noticed to occur within NLO pQCD [39], where the introduction of the strange quark mass was observed to push down the pressure at low-$\mu_{B}$ values ($\mu_{B}\lesssim 3$ GeV), softening the resulting EoS. However, later the authors of Ref. [35] found that the softening effect is less noticeable when finite masses are considered at NNLO.
[…]

### v2

@@ Abstract
We employ the renormalization group optimized perturbation theory (RGOPT) resummation method to evaluate the equation of state (EoS) for strange ($N_{f}=2+1$) and non-strange ($N_{f}=2$) cold quark matter at next-to-leading order (NLO). This allows us to obtain the mass-radius relation for pure quark stars and compare the results with the predictions from perturbative QCD (pQCD) at NNLO. Choosing the renormalization scale to generate maximum star masses of order $M=2-2.6M_{\odot}$, we show that the RGOPT can produce mass-radius curves compatible with the masses and radii of some recently observed pulsars, regardless of their strangeness content. The scale values required to produce the desired maximum masses are higher in the strange scenario since the EoS is softer in this case. The possible reasons for such behavior are discussed. Our results also show that, as expected, the RGOPT predictions for the relevant observables are less sensitive to scale variations than those furnished by pQCD.
@@ I Introduction
Nevertheless, it would be legitimate to argue that such a result could be an artifact of the model approximation adopted in Ref. [20]. After all, when compared to QCD, the (effective) NJL model has some limitations which include the absence of asymptotic freedom and non-renormalizability. On the other hand, the determination of the QCD EoS for cold and compressed matter, from ab initio evaluations, presents further issues which have not yet been circumvented. For instance, within this regime lattice QCD (LQCD) simulations are still plagued by the infamous sign problem [33, 34] while perturbative QCD (pQCD) applications are only reliable at extremely high baryon densities, of order $n_{B}\sim 40n_{0}$ [35, 36, 37, 38] ($n_{0}=0.16\,{\rm fm}^{-3}$), where asymptotic freedom allows for weak coupling expansions. Regarding pQCD it is important to mention that when the modified minimal subtraction ($\overline{\rm MS}$) renormalization scale ($\Lambda$) is taken at the conventional “central” value, $2(\mu_{u}+\mu_{d}+\mu_{s})/3$, with $\mu_{f}$ the quark flavor chemical potentials, the next-to-leading order NLO pQCD predicts QS masses above $2M_{\odot}$ [39] and below $2M_{\odot}$ at NNLO [40, 41, 35, 42, 43].
@@
Although the cold dense pQCD NNLO pressure is formally perturbatively renormalization group (RG) invariant, the higher order residual scale dependence of pQCD results [35, 36, 37] is still rather sizeable, conventionally taken to lie between half and twice the central scale choice value. Quite recently, pQCD calculations reached NNNLO [44, 45, 46, 47] within the massless quark approximation, and all orders in the leading soft logarithms approximation [48], producing in such cases significant residual scale dependence reductions. However, accounting for the strange quark mass effects, sizeable at moderate and low $\mu$ values, is very involved and presently limited to NNLO, thus impiging on scale uncertainties reductions. In order to tackle the scale dependence problem an alternative method, which better incorporates RG properties within (variationally) optimized perturbative evaluations has been proposed [49]. This renormalization group optimized perturbation theory (RGOPT) technique has been successfully used, at vanishing temperatures and densities, to calculate quantities such as the QCD scale $\Lambda_{\overline{\rm MS}}$ and strong coupling $\alpha_{s}$ [49], and the quark condensate up to four and five loop orders [50, 51]. Later, the RGOPT has been applied to scalar field theories at finite temperatures [52, 53, 54, 55], indicating that the method greatly improves the scale dependence in thermal theories, when compared to screened perturbation theory (SPT) [56, 57] and hard thermal loop perturbation theory (HTLpt) [58, 57, 59] based on the HTL effective theory [60]. The technique was also employed to describe cold and dense QCD [61], as well as the quark sector of QCD at finite temperatures and baryonic densities [62, 63]. Once again, the RGOPT predictions display a residual scale dependence which is smaller than those furnished by pQCD and HTLpt. Recently, some of the present authors have used the RGOPT EoS for QCD in the chiral limit ($m_{u}=m_{d}\equiv 0$) to describe non-strange quark stars (NSQS) obtaining encouraging results at NLO [32]. In particular, when the renormalization scale is appropriately chosen, the RGOPT resummation of the perturbative series is able to describe several observational data, specially from pulsars PSR J0030+0451 [64, 65], PSR J0740+6620 [66, 67, 68] and the light compact object identified in the middle of a supernova remnant, HESS J1731-347 [69]. Note that, recently a different approach has been considered to constrain the scale $X$ within a model, possessing a bag parameter constrained by LQCD, which describes an isospin-dense matter system [70].
@@
Here, our major goal is to investigate how the inclusion of strangeness affects the $N_{f}=2$ results obtained in Ref. [32]. In the present $N_{f}=2+1$ case, the introduction of a finite (strange) quark mass ($m_{s}$) modifies the RG operator introducing an extra scale dependence, through the running of $m_{s}(\Lambda)$, which needs to be taken into account by the resummation procedure. This is accordingly a nontrivial extension with respect to the determination of the RGOPT EoS for massless quarks at $N_{f}=3$ [61, 62, 63] and $N_{f}=2$ [32]. Following the usual procedure, here we will also choose the scale to be density dependent so that it can be written in terms of the baryonic chemical potential [^1] In this situation the pressure needs to be modified to ensure thermodynamic consistency. ($\mu_{B}$) as $\Lambda=X\mu_{B}/3$, where $X$ represents a free parameter that can be adjusted in order to reproduce the desired maximum mass value for a given family of stars. Then, the QS properties determined from the novel $N_{f}=2+1$ RGOPT EoS will be compared with those previously obtained with the same method when $N_{f}=2$ [32] as well as with those furnished by $N_{f}=2+1$ pQCD at NNLO [35, 36], for the corresponding flavor species. We stress that our approach is not a fixed order perturbative calculation but actually resumming a certain class of higher order contributions, indeed in the massless quark approximation the NLO RGOPT pressure is definitely numerically closer to NNLO pQCD than to NLO pQCD [61], and the latter would be a poor approximation to consider compared to the state-of-the art NNLO pQCD. Therefore, in the present case with $m_{s}\neq 0$ we also consider appropriate to compare NLO RGOPT with NNLO pQCD. Since $\alpha_{s}$ runs with $\Lambda=X\mu_{B}/3$ the value of $X$ needed to produce a massive quark star will enable us to compare the coupling strength associated with each different number of flavors. This comparison will then allow us to draw further conclusions about the reliability of the distinct EoS provided by the RGOPT at NLO.
@@ II Perturbative QCD pressure for $N_{f}=2+1$ at NLO
[…]
@@ III RGOPT resummation setup
At this point, some important additional remarks are in order: First, note that although these RG properties were originally obtained from vacuum ($\mu=T=0$) contributions, they still hold when thermal and/or in-medium effects are considered [53, 52, 54, 62, 63]. Also, according to the original prescription, $a$ is fixed once for all at LO, and the same value in Eq. (15) is used at successive perturbative orders of the (modified) perturbative $\delta$-expansion. Importantly, this guarantees [49] in particular that at higher orders there is always one solution $\overline{\eta}(g)$ compatible with asymptotic freedom behavior for $g\to 0$. Finally, note that even when treating massive theories as we do here, it is compelling to adopt this universal constant value for $a$ in Eq. (15), in order to avoid otherwise a clumsy a(g,m,)a(g,m,...) dependent exponent, furthermore incompatible with standard renormalization. We emphasize that although other slightly different variational prescriptions could be thought of, the latter one provides a sensible way of comparing successive perturbative orders with the same prescription. This is further justified a posteriori since the strange quark mass ($m_{s}$) which is relevant here remains a rather moderate perturbation [^4] As we will see in the sequel, $m_{s}\ll\overline{\eta}\simeq{\cal O}(g\,\mu)\ll\mu$ roughly holds, at least as long as the QCD coupling, $g^{2}(\Lambda)$, remains moderate (i.e. for renormalization scales $\Lambda\sim{\cal O}(\mu)$ with $\mu$ not too small). with respect to the dressed mass $\eta(g^{2},\mu)$ determined from Eq. (17).
@@ III.1 Optimization
After applying this RSC to the original NLO pressure, one must re-expand perturbatively the latter to NLO, producing an extra term $4g^{2}\,m_{f}^{4}s_{0}B_{2}$ in the resulting pressure. As a by-product, after the RGOPT modifications, similar contributions with $m_{f}\to\eta+m_{s}$ and $m_{f}\to\eta$ are induced respectively in the first and second term on the right hand side of Eq. (21). Although the dimensionless RSC parameter $B_{2}$ in Eq.(23) is arbitrary to start with, one seeks a sensible prescription to recover real solutions while maintaining $B_{2}$ reasonably close to $\overline{\rm MS}$ results to justify a posteriori such a perturbative RSC. As Ref. [49] prescribes, $B_{2}$ is accordingly fixed (uniquely) upon requiring the nearest “contact” of the two curves parametrizing the MOP Eq. (17) and RG Eq. (16), considered as functions $f_{\rm{MOP}}(\eta,g^{2})$, $f_{\rm{RG}}(\eta,g^{2})$ respectively. Algebraically this corresponds to solving
@@
Fig. 1 compares the optimized mass parameter, $\overline{\eta}$, as well as the normalized pressure as functions of the baryon chemical potential representing symmetric matter, $\mu_{B}=3\mu$, for $N_{f}=3$ and $N_{f}=2+1$. Following the standard prescription, we have considered three different values of the renormalization scale, $\Lambda=X\mu_{B}/3$, with $X=1,2,4$, to quantify the residual scale dependence. Regarding the variational masses, the figure suggests that they share a similar behavior, increasing linearly with $\mu_{B}$ roughly as $\overline{\eta}\sim{\cal O}(g\,\mu)$, although $\overline{\eta}$ actually entails an all-order $g^{2}$-dependence from RG resummation. Importantly, the fact that $\overline{\eta}$ accordingly behaves as a screening mass for moderate coupling $g$, also guarantees that the RSC Eq. (23) modifications induced to the pressure, $\propto-B_{2}s_{0}g^{2}\overline{\eta}^{4}$, are of formally higher orders ${\cal O}(g^{6})$, so that the physical pressure is only mildly dependent on the renormalization scheme choice. For the same $\mu_{B}$ value, and comparing with both $N_{f}=2$ and $N_{f}=3$ cases, the presence of a massive quark ($N_{f}=2+1$) generates a lower pressure. This can be explained by first noting that in the basic NLO pressure Eq. (1) or (2), the $m_{f}\neq 0$ contributions are always negative as compared to the equivalent massless contributions (at least as long as $m_{f}\ll\mu$), and this feature remains true for the modified NLO expression Eq. (19). Then since $\overline{\eta}+m_{s}>\overline{\eta}$, the lowering of the pressure is obviously more pronounced for $m_{s}\neq 0$ in the RGOPT case [^7] In contrast, the RSC modification $\sim-B_{2}s_{0}g^{2}\overline{\eta}^{4}$ gives a positive extra contribution to the pressure, as $s_{0}>0$ in Eq. (6) and we always obtain $B_{2}<0$. But this contribution is subdominant compared to the ”direct” $\overline{\eta}$ dependence entering other contributions.. A lower pressure tends to enhance the remnant scale dependence, therefore slightly counteracting the scale dependence improvement obtained from RGOPT, although the figure indicates that the observed scale dependence is not so much sensitive to the strange quark mass value.
@@ IV Chemical equilibrium and thermodynamic consistency
must be observed. The relevant baryon chemical potential describing each different number of flavors is defined as $\mu_{B}=\mu_{u}+2\mu_{d}$ (recall that $\mu_{d}=\mu_{s}$). Seeking a fair comparison, the renormalization scale given by $\Lambda=X\,\mu_{B}/3$ is chosen in all cases.
At this stage we should specify how to best compare our RGOPT pressure results with pQCD, which for the full $m_{s}$ dependence is known at NNLO [35, 36, 78]. As anticipated in the introduction we consider more appropriate to compare NLO RGOPT with NNLO pQCD, since the NLO RGOPT expression given by Eq.(19) resums an all order RG dependence, and in the massless quark approximation the NLO RGOPT pressure is numerically closer [61] to NNLO pQCD than to NLO pQCD (we have also explicitly checked that a similar behavior holds for $m_{s}\neq 0$). To obtain the pQCD results at $N_{f}=2$, we thus use the NNLO result for massless quarks given by [40, 41, 79],
@@ V Numerical Results
In this section, we first present the numerical results furnished by the RGOPT and pQCD for NSQM ($N_{f}=2$) as well as SQM ($N_{f}=2+1$) considering $\beta$-equilibrium, Eqs. (27) and (29), and charge neutrality, Eqs. (30) and (28). Next, the EoS corresponding to each case is used to obtain the mass-radii relation describing non-strange and strange quark stars.
V.1 NSQM and SQM for $\beta$-equilibrium matter
@@
The electron chemical potential predicted by the NLO RGOPT is plotted in Fig. 2 (on the left panel) as a function of $\mu_{B}$ for $N_{f}=2+1$ (solid lines) and $N_{f}=2$ (dashed lines). In the case of SQM, the appearance of kinks in the curves signals the onset of strange quarks which induces a strong reduction of the number of electrons decreasing the electron chemical potential ($\mu_{e}\to 0$, at high-$\mu_{B}$). Therefore, in this limit, the densities of the three quark flavors become similar. In other words, the system favors the replacement of the negative electric charge of electrons by the one of $s$-quarks. On the other hand, in the case of NSQM, as $\mu_{B}$ takes on higher values, the electron production increases to compensate for the increase in the $u$ quark number density, while maintaining charge neutrality. Note that the $u$ quark has a smaller Fermi momentum than the $d$ quark and, therefore, an increase of the number density is favored.
@@
Figure 3: Thermodynamic consistent NLO RGOPT pressure as function of $\mu_{B}$ for $N_{f}=2+1$ (continuous lines) and $N_{f}=2$ (dashed lines). For comparison the dot-dashed and dotted lines respectively represent the $N_{f}=2+1$ and $_{f}=2 NN$LO pQCD pressure. For all curves the upper ones correspond to $X=4$, the lower ones to $X=1$ while the central scale, $X=2$, is represented by the inner curves. All pressures have been normalized by the Fermi-Dirac value $P_{\rm FD}$ for free massless quarks.
@@
In Fig. 3, we show the normalized pressures furnished by the NLO RGOPT and NNLO pQCD for both SQM and NSQM, thermodynamically consistent after applying Eq. (34). The first thing to note is that, within both approximations, the introduction of a finite strange quark mass decreases the scale dependence (this effect is much more noticeable in the case of pQCD). While those features result from involved combined effects from the running coupling, running mass, and extra thermodynamic consistency contributions, this behavior is explained partly from the fact that since $b_{0}(N_{f}=3)<b_{0}(N_{f}=2)$, the running coupling $g^{2}(\Lambda)\sim[2b_{0}\ln(\Lambda/\Lambda_{QCD})\,]^{-1}$ has somewhat smaller scale variation for $N_{f}=3$ than for $N_{f}=2$. The latter effect is however partly compensated by the sizable massive contributions $\overline{\eta}\neq 0$, lowering the pressure (as explained in the last paragraph before Sec. III), which rather tends to increase the remnant scale dependence. As explained in Sec. III, the RGOPT procedure involves a resummed all order RG dependence, thus adding higher order $\Lambda$-dependence that can further reduce the renormalization scale dependence, so that our results display a somewhat reduced scale dependence with respect to unresummed pQCD. However, the improvement that we obtain for the present cold quark matter case is much more moderate than its counterpart for hot QCD in [62, 63]. In addition, the figure suggests that the NLO RGOPT produces EoSs that converge slower to the Fermi-Dirac limit (free massless quarks) than its NNLO pQCD counterpart. This happens because the former procedure produces a non-trivial medium-dressed mass $\overline{\eta}(g,\mu)$, so that the effective masses entering the quark propagators are $m_{s}+\overline{\eta}(g,\mu)$ (strange sector) and $\overline{\eta}(g,\mu)$ (non-strange sector). Accordingly, at $N_{f}=2$ the RGOPT quark propagator already involves a non-trivial effective quark mass while, at $N_{f}=2+1$, the effect of including the $s$-quark physical (current) mass is more diluted. A similar behavior was initially noticed to occur within NLO pQCD [39], where the introduction of the strange quark mass was observed to push down the pressure at low-$\mu_{B}$ values ($\mu_{B}\lesssim 3$ GeV), softening the resulting EoS. However, later the authors of Ref. [35] found that the softening effect is less noticeable when finite masses are considered at NNLO.
V.2 Quark stars
Figure 4: Pressure versus baryon chemical potential (top left) and energy density (top right) of $\beta$-equilibrated matter for the $N_{f}=2+1$ (continuous lines) and $N_{f}=2$ (dashed lines) NLO RGOPT, and for $N_{f}=2+1$ (dot-dashed lines) and $N_{f}=2$ (dotted lines) NNLO pQCD. Bottom panel: the quark fractions as a function of the baryonic density for the same EoS. The curves were obtained with the scale values $X$ that reproduce $M_{\rm max}=2,\,2.3$ and $2.6M_{\odot}$, given in Table 1. The dots identify the values at the center of the maximum mass star configuration. The squares in the bottom panel identify the baryonic densities at which $P=0$.